OpenStudy (anonymous):

I need some help! Simplify. 4 times the 4th root of 128

OpenStudy (anonymous):

\[\large 4 \times \sqrt[4]{128} = 4 \cdot \sqrt[4]{2^4 \cdot 2^3}\]

OpenStudy (anonymous):

Now you can take out one 2 out of brackets: \[\large \implies 4 \cdot 2 \sqrt[4]{2^3}= ??\]

OpenStudy (anonymous):

so its \[8\sqrt[4]{8} \]

OpenStudy (anonymous):

Yep...

OpenStudy (anonymous):

ok thank you!

OpenStudy (anonymous):

You are welcome dear...